# Intuitionistic NF set theory

### v5

@@ 3. Finite sets
The set FINITE of finite sets is defined as the intersection of all $X$ such that $X$ contains the empty set $\Lambda$ and
@@
The formula in the definition can be stratified by giving $u$ index 1, $z$ index 0, and $X$ index 2, so the definition can be given in INF.
@@
Proof. Let $Z$ be the set of finite sets with decidable equality. I say that $Z$ satisfies the closure conditions in the definition of FINITE, Definition 3.1. The empty set has decidable equality, so the first condition holds. Now suppose $Y=X\cup\{a\}$, where $X\in Z$ and $a\not\in X$. We must show $Y\in Z$. Let $x,y\in Y$. Then $x\in X\lor x=a$ and $y\in X\lor y=a$. There are thus four cases to consider: If both $x$ and $y$ are in $X$, then by the induction hypothesis, we have the desired $x=y\lor x\neq y$. If one of $x,y$ is in $X$ and the other is $a$, then $x\neq y$, since $a\not\in X$; hence $x=y\lor x\neq y$. Finally if both are equal to $a$, then $x=y$ and hence $x=y\lor x\neq y$. Therefore, as claimed, $Z$ satisfies the closure conditions. Hence every finite set belongs to $Z$. That completes the proof of the lemma.
@@
Proof. Define
$Z=\{X\in{\sf FINITE}\ :X=\Lambda\ \lor\ \exists u\,(u\in X)\}.$
We will show $Z$ satisfies the closure conditions in the definition of FINITE. Evidently $\Lambda\in Z$. Now suppose $X\in Z$ and $Y=X\cup\{a\}$ with $a\not\in X$. We must show $Y\in Z$. Since $X\in Z$, $X$ is finite. Therefore $Y$ is finite. Since $a\in Y$ we have $Y\in Z$. That completes the proof of the lemma.
@@
Proof. Let $X$ be a finite set. Suppose $\neg\neg\,\exists u\,(u\in X)$. That is, $X$ is nonempty. By Lemma 3.4, $X$ has a member. That completes the proof of the lemma.
@@
$\Lambda\in{\sf FINITE}$.
Proof. $\Lambda$ belongs to every set $W$ containing $\Lambda$ and containing $u\cup\{e\}$ whenever $u\in W$ and $e\not\in W$. Since ${\sf FINITE}$ is the intersection of such sets $W$, $\Lambda\in{\sf FINITE}$. That completes the proof of the lemma.
@@
Proof. Let $x\in{\sf FINITE}$. Then $x$ belongs to every set $W$ containing $\Lambda$ and containing $u\cup\{e\}$ whenever $u\in W$ and $e\not\in W$. Let $W$ be any such set. Then $x\cup\{c\}\in W$. Since $W$ was arbitrary, $x\cup\{c\}\in{\sf FINITE}$. That completes the proof of the lemma.
@@
If $z\in{\sf FINITE}$, then $z=\Lambda$ or there exist $x\in{\sf FINITE}$ and $c\not\in x$ such that $z=x\cup\{c\}$.
Proof. The formula is stratified, giving $c$ index 0, and $x$ and $z$ index 1. ${\sf FINITE}$ is a parameter. We prove the formula by induction on finite sets. Both the base case (when $z=\Lambda$) and the induction step are immediate. That completes the proof of the lemma.
@@
Proof. We have
$\Lambda\in{\sf FINITE}$   by Lemma 3.6
$x\not\in\Lambda$   by the definition of $\Lambda$
$\Lambda\cup\{x\}\in{\sf FINITE}$   by Lemma 3.7
$\{x\}=\Lambda\cup\{x\}$   by the definitions of $\cup$ and $\Lambda$
@@
$USC(x)$ is finite if and only if $x$ is finite.
Proof. Left-to-right: we have to prove
(2) $\forall y\in{\sf FINITE}\,\forall x\,(y=USC(x)\ {\rightarrow}\ x\in{\sf FINITE})$
@@
Base case. When $y=\Lambda=USC(x)$ we have $x=\Lambda$, so $x\in{\sf FINITE}$.
Induction step. Suppose $y\in{\sf FINITE}$ has the form $y=z\cup\{w\}=USC(x)$ and $w\not\in z$, and $z\in{\sf FINITE}$. Then $w=\{c\}$ for some $c\in x$. The induction hypothesis is
(3) $\forall w\,(z=USC(w)\ {\rightarrow}\ w\in{\sf FINITE})$
@@
$=$ $USC(x)-\{\{c\}\}\mbox{\qquad\ since $y=USC(x)$ and $w=\{c\}$}$
$=$ $USC(x-\{c\}).$
@@
(5) $x\in{\sf FINITE}\ {\rightarrow}\ USC(x)\in{\sf FINITE}$
@@
Base case: $USC(\Lambda)=\Lambda\in{\sf FINITE}$.
@@
$USC(x\cup\{c\})=USC(x)\cup\{\{c\}\}.$
Let $c\not\in x$ and $x\in{\sf FINITE}$. By the induction hypothesis (5), $USC(x)$ is finite, and since $c\not\in x$, we have $\{c\}\not\in USC(x)$. Then $USC(x)\cup\{\{c\}\}$ is finite. Then $USC(x\cup\{c\})$ is finite. That completes the induction step. That completes the proof of the lemma.
@@
Proof. We prove by induction on finite sets $X$ that
$\forall\,Y\in{\sf FINITE}\,(X\cap Y=\Lambda\ {\rightarrow}\ X\cup Y\in{\sf FINITE}).$
Base case: $\Lambda\cup Y=Y$ is finite.
Induction step: Suppose $X=Z\cup\{b\}$ with $b\not\in Z$ and $Y\cap(Z\cup\{b\})=\Lambda$ and $Z$ finite. Then
@@
(6) $X\cup Y$ $=$ $Z\cup(Y\cup\{b\})$
Since $Y\cap(Z\cup\{b\})=\Lambda$, $b\not\in Y$. Then by the definition of ${\sf FINITE}$, $Y\cup\{b\}$ is finite. We have
$Z\cap(Y\cup\{b\})=Y\cap(Z\cup\{b\})=\Lambda.$
Then by the induction hypothesis, $Z\cup(Y\cup\{b\})$ is finite. Then by (6), $X\cup Y$ is finite. That completes the induction step. That completes the proof of the lemma.
@@
Proof. Suppose $x\sim y$. Then there exists $f:x\to y$ with $f$ one-to-one and onto. By Lemma 2.9, $f^{-1}:y\to x$ is a one-to-one function. Then we have for $u,v\in y$,
(7) $u=v{\ \leftrightarrow\ }f^{1}(u)=f^{-1}(v).$
@@
Therefore $y$ has decidable equality.
@@
Proof. Let $q=f(c)$. Then $g:z\to y-\{q\}$. Suppose $g(u)=g(v)$. Then $f(u)=f(v)$. Since $f$ is one-to-one, $u=v$. Hence $g$ is one-to-one. Suppose $v\in y-\{q\}$. Since $f$ is onto, $v=f(u)$ for some $u\in z\cup\{c\}$; but $u\neq c$ since if $u=c$ then $v=f(u)=q$, but $v\neq q$ since $v\in y-\{q\}$. Then $u\in z$. Hence $g$ is onto. That completes the proof.
@@
Proof. We prove by induction on finite sets $x$ that
@@
Base case: When $x=\Lambda$. Suppose $y\sim\Lambda$. Then $y=\Lambda$, so $y\in{\sf FINITE}$. That completes the base case.
@@
Therefore $y\in{\sf FINITE}$. That completes the induction step. That completes the proof of the lemma.
@@
$SC(X)=\{Y:Y\subset X\}.$
We shall not make use of $SC(X)$, because there are “too many” subclasses of $X$. Consider, by contrast, the separable subclasses of $X$:
@@
$SSC(X):=\{u:u\subseteq X\ \land\ X=u\cup(X-u)\}$
That is, $u$ is a separable subclass (or subset, which is synonymous) of $X$ if and only if $\forall y\in X\,(y\in u\ \lor\ y\not\in u)$. Classically, of course, every subset is separable, so we have $SSC(X)=SC(X)$, but that is not something we can assert constructively. The formula in the definition is stratified, so the definition can be given in INF. When working with finite sets, $SSC(X)$ is a good constructive substitute for $SC(X)$. We illustrate this by proving some facts about $SC(X)$, before returning to the question of the proper constructive substitute for $SC(X)$ when $X$ is not necessarily finite.
@@
Let $x$ be a finite set. Then $SSC(x)$ is also a finite set.
Remark. We cannot prove this with $SC(x)$ in place of $SSC(x)$.
Proof. The formula to be proved is
$x\in{\sf FINITE}\ {\rightarrow}\ SSC(x)\in{\sf FINITE}.$
@@
Base case: $SSC(\Lambda)=\{\Lambda\}$ is finite.
@@
By the induction hypothesis, $SSC(x)\in{\sf FINITE}$. I say that the map $u\mapsto u\cup\{c\}$ is definable in INF:
$f:=\{\langle u,y\rangle:u\in SSC(x)\ \land\ y=u\cup\{c\}.\}$
The formula can be stratified by giving $c$ index 0, $u$ and $y$ index 1, $SSC(x)$ index 2; then $\langle u,y\rangle$ has index 3 and we can give $f$ index 4. Hence $f$ is definable in INF as claimed. $f$ is a function since $y$ is uniquely determined as $u\cup\{c\}$ when $u$ is given. Also $f$ is one-to-one, since if $u\subseteq x$ and $v\subseteq x$ and $c\not\in x$, and $u\cup\{c\}=v\cup\{c\}$, then $u=v$. Define
@@
(8) $A=\{u\cup\{c\}:u\in SSC(x)\}.$
Then $SSC(x)\sim A$, because $f:SSC(x)\to A$ is one-to-one and onto. Since $SSC(x)$ is finite (by the induction hypothesis), by Lemma 3.3, $SSC(x)$ has decidable equality. Then $A$ has decidable equality, by Lemma 3.12. Since $A$ has decidable equality, and is similar to the finite set $SSC(x)$, $A$ is finite, by Lemma 3.14.
@@
Left-to-right: Let $v\in SSC(x\cup\{c\})$. Then $v$ is a separable subset of $x\cup\{c\}$. Then $c\in v\ \lor\ c\not\in v$. If $c\not\in v$ then $v\in SSC(x)$. If $c\in v$
@@
We have $v-\{c\}\in SSC(x)$, since $v\subset x\cup\{c\}$ and $v$ has decidable equality. Then $f(v-\{c\})\in Range(f)=A$. But $f(v-\{c\})=(v-\{c\})\cup\{c\}=v$. Therefore $v\in A$. Therefore $v\in A\cup SSC(x)$, as desired. That completes the proof of the left-to-right direction of (9).
Right-to-left. Let $v\in A\cup SSC(x)$. Then $v\in A\ \lor\ v\in SSC(x)$.
Case 1, $v\in A$. Then by (8), $v$ has the form $v=u\cup\{c\}$ for some $u\in SSC(x)$. Then $u\cup\{c\}\in SSC(x\cup\{c\}$ as required.
Case 2, $v\in SSC(x)$. First we note that if $c\not\in x$ then
$SSC(x)\subseteq SSC(x\cup\{c\})$
Therefore, since $v\in SSC(x)$, we have $v\in SSC(x\cup\{c\})$. That completes the proof of (9).
Note that $A$ and $SSC(x)$ are disjoint, since every member of $A$ contains $c$, and no member of $SSC(x)$ contains $c$, since $c\not\in x$. Then by Lemma 3.11 and (9), $SSC(x\cup\{c\})\in{\sf FINITE}$, as desired. That completes the proof of the lemma.
@@
Proof. Let $a\in{\sf FINITE}$. By induction on finite sets $b$ we prove
(10) $b\in{\sf FINITE}\ {\rightarrow}\ b\subseteq a\ {\rightarrow}\ a=(a-b)\cup b.$
@@
Base case: Suppose $b=\Lambda$. Then $b\subseteq a$, so we have to prove $a=(a-\Lambda)\cup\Lambda$, which is immediate. That completes the base case.
@@
We must show $x\in a$. If $x\in(a-(b\cup\{c\})$ then $x\in a$. If $x\in(b\cup\{c\})$ then $x\in a$, since by hypothesis $b\cup\{c\}\subseteq a$. That completes the right-to-left direction. That completes the induction step. That completes the proof of the lemma.
@@
Proof. By induction on finite sets $X$. When $X$ is the empty set, every subset of $X$ is the empty set, so every subset of $X$ is empty, and hence finite. Now let $X=Y\cup\{a\}$ with $a\not\in Y$ and $Y$ finite, and let $U$ be a separable subset of $X$; that is,
(15) $\forall z\in X\,(z\in U\ \lor\ z\not\in U).$
@@
(16) $\forall z\in Y\,(z\in V\ \lor\ z\not\in V)$
Let $z\in Y$. Since $U$ is a separable subset of $X$, $z\in U\lor z\not\in U$. By Lemma 3.3, $X$ has decidable equality, so $z=a\lor z\neq a$. Therefore $z\in V\lor z\not\in V$, as claimed in (16). Then, by the induction hypothesis, $V$ is finite. Since $a\not\in V$, also $V\cup\{a\}$ is finite. I say that $V\cup\{a\}=U$. If $x\in V\cup\{a\}$ then $x\in U$, since $V\subseteq U$ and $a\in U$. Conversely if $x\in U$ then $x=a\ \lor x\neq a$, since $a$ and $x$ both are members of $X$ and $X$ has decidable equality by Lemma 3.3. If $x=a$ then $x\in\{a\}$ and if $x\neq a$ then $x\in V$, so in either case $x\in V\cup\{a\}$. Therefore $V\cup\{a\}=U$ as claimed. Since $V$ is finite and $a\not\in V$, $V\cup\{a\}$ is finite. Since $U=V\cup\{a\}$, $U$ is finite. That completes the induction step. That completes the proof of the lemma.
@@
Let $a$ and $b$ be finite sets with $b\subseteq a$. Then $a-b$ is also a finite set.
Proof. We first prove the special case when $b$ is a singleton, $b=\{c\}$. That is,
(17) $a\in{\sf FINITE}\ \land\ c\in a\ {\rightarrow}\ a-\{c\}\in{\sf FINITE}$
By Lemma 3.3, $a$ has decidable equality. Hence $a-\{c\}$ is a separable subset of $a$. Then by Lemma 3.19, it is finite. That completes the proof of 17.
@@
Base case: $\Lambda-b=\Lambda$ is finite.
@@
Since $b$ is finite, also $b-\{c\}$ is finite, by 17. Since $b-\{c\}\subseteq p$, by the induction hypothesis we have
@@
Case 2: $c\not\in b$. Then $b\subseteq p$, so by the induction hypothesis $p-b$ is finite.
@@
Therefore $a-b$ is finite. That completes Case 2. That completes the induction step. That completes the proof of the lemma.
@@
Proof. The formula to be proved is stratified, with ${\sf FINITE}$ as a parameter, giving $u$ and $z$ index 0, $B$ index 1, and $R$ index 3. Therefore it is legal to prove it by induction on finite sets $B$.
Base case: $B=\Lambda$. Then $X\times Y=\Lambda$, so $R=\Lambda$. Then $\forall z\,\neg\,\exists u\in B\,\langle u,z\rangle\in R$, and therefore
@@
[…]
@@
(19) $(\exists u\in B\,\langle u,z\rangle\in R)\ \lor\ \neg\,(\exists u\in B\,\langle u,z\rangle\in R)$
@@
[…]
@@
But by the induction hypothesis, that holds. That completes the induction step. That completes the proof of the lemma.
@@
Proof. Since $b\in U$ and $c\not\in U$, we have $b\neq c$. Define $f:U\to Y$ by
@@
[…]
@@
The class $X$ is infinite if $X\sim Y$ for some $Y\subseteq X$ with $Y\neq X$.
@@
Proof. It suffices to show that every finite set is not infinite. The formula to be proved is

### v6

@@ 3 Finite sets
The set FINITE of finite sets is defined as the intersection of all $X$ such that $X$ contains the empty set $\emptyset$ and
@@
The formula in the definition can be stratified by giving $u$ index 1, $z$ index 0, and $X$ index 2, so the definition can be given in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
This definition was introduced in [5] as “$N$-finite.” [^9] He also defined other notions of “finite”; for example $K$-finite drops the requirement $z\not\in u$ from the definition. That notion, and the other notion considered op. cit., do not satisfy the property that the cardinality of a finite set is a finite cardinal, i.e., an integer. For example, $\{c\}$ will be $K$-finite, even if we do not know whether or not $c$ is inhabited, so we cannot assign $\{c\}$ a finite cardinal.
@@
Proof.
Let $Z$ be the set of finite sets with decidable equality. I say that $Z$ satisfies the closure conditions in the definition of FINITE, Definition 3.1. The empty set has decidable equality, so the first condition holds. Now suppose $Y=X\cup\{a\}$, where $X\in Z$ and $a\not\in X$. We must show $Y\in Z$. Let $x,y\in Y$. Then $x\in X\lor x=a$ and $y\in X\lor y=a$. There are thus four cases to consider: If both $x$ and $y$ are in $X$, then by the induction hypothesis, we have the desired $x=y\lor x\neq y$. If one of $x,y$ is in $X$ and the other is $a$, then $x\neq y$, since $a\not\in X$; hence $x=y\lor x\neq y$. Finally if both are equal to $a$, then $x=y$ and hence $x=y\lor x\neq y$. Therefore, as claimed, $Z$ satisfies the closure conditions. Hence every finite set belongs to $Z$. ∎
@@
Proof.
Define
$Z=\{X\in{\sf FINITE}\ :X=\emptyset\ \lor\ \exists u\,(u\in X)\}.$
We will show $Z$ satisfies the closure conditions in the definition of FINITE. Evidently $\emptyset\in Z$. Now suppose $X\in Z$ and $Y=X\cup\{a\}$ with $a\not\in X$. We must show $Y\in Z$. Since $\in Z, XX$ is finite. Therefore $Y$ is finite. Since $a\in Y$ we have $Y\in Z$. ∎
@@
Proof.
Let $X$ be a finite set. Suppose $\neg\neg\,\exists u\,(u\in X)$. That is, $X$ is nonempty. By Lemma 3.4, $X$ has a member. ∎
@@
$\emptyset\in{\sf FINITE}$.
Proof.
$\emptyset$ belongs to every set $W$ containing $\emptyset$ and containing $u\cup\{e\}$ whenever $u\in W$ and $e\not\in W$. Since $\sf FINITE$} is the intersection of such sets $W$, $\emptyset\in{\sf FINITE}$. ∎
@@
Proof.
Let $x\in{\sf FINITE}$. Then $x$ belongs to every set $W$ containing $\emptyset$ and containing $u\cup\{e\}$ whenever $u\in W$ and $e\not\in W$. Let $W$ be any such set. Then $x\cup\{c\}\in W$. Since $W$ was arbitrary, $x\cup\{c\}\in{\sf FINITE}$. ∎
@@
If $z\in{\sf FINITE}$, then $z=\emptyset$ or there exist $x\in{\sf FINITE}$ and $c\not\in x$ such that $z=x\cup\{c\}$.
Proof.
The formula is stratified, giving $c$ index 0, and $x$ and $z$ index 1. $\sf FINITE$} is a parameter. We prove the formula by induction on finite sets. Both the base case (when $z=\emptyset$) and the induction step are immediate. ∎
@@
Proof.
We have
$\emptyset\in{\sf FINITE}$   by Lemma 3.6
$x\not\in\emptyset$   by the definition of $\emptyset$
$\emptyset\cup\{x\}\in{\sf FINITE}$   by Lemma 3.7
$\{x\}=\emptyset\cup\{x\}$   by the definitions of $\cup$ and $\emptyset$
@@
$\mathcal{P}_{1}(x)$ is finite if and only if $x$ is finite.
Proof.
Left-to-right: we have to prove
$\forall y\in{\sf FINITE}\,\forall x\,(y=\mathcal{P}_{1}(x)\ {\rightarrow}\ x\in{\sf FINITE})$ (2)
@@
Base case. When $y=\emptyset=\mathcal{P}_{1}(x)$ we have $x=\emptyset$, so $x\in{\sf FINITE}$.
Induction step. Suppose $y\in{\sf FINITE}$ has the form $y=z\cup\{w\}=\mathcal{P}_{1}(x)$ and $w\not\in z$, and $z\in{\sf FINITE}$. Then $w=\{c\}$ for some $c\in x$. The induction hypothesis is
$\forall w\,(z=\mathcal{P}_{1}(w)\ {\rightarrow}\ w\in{\sf FINITE})$ (3)
@@
$=$ $\mathcal{P}_{1}(x)-\{\{c\}\}\mbox{\qquad\ since $y=\mathcal{P}_{1}(x)$ and $w=\{c\}$}$
$=$ $\mathcal{P}_{1}(x-\{c\}).$
@@
$x\in{\sf FINITE}\ {\rightarrow}\ \mathcal{P}_{1}(x)\in{\sf FINITE}$ (5)
@@
Base case: $\mathcal{P}_{1}(\emptyset)=\emptyset\in{\sf FINITE}$.
@@
$\mathcal{P}_{1}(x\cup\{c\})=\mathcal{P}_{1}(x)\cup\{\{c\}\}.$
Let $c\not\in x$ and $x\in{\sf FINITE}$. By the induction hypothesis (5), $\mathcal{P}_{1}(x)$ is finite, and since $c\not\in x$, we have $\{c\}\not\in\mathcal{P}_{1}(x)$. Then $\mathcal{P}_{1}(x)\cup\{\{c\}\}$ is finite. Then $\mathcal{P}_{1}(x\cup\{c\})$ is finite. That completes the induction step. ∎
@@
Proof.
We prove by induction on finite sets $X$ that
$\forall\,Y\in{\sf FINITE}\,(X\cap Y=\emptyset\ {\rightarrow}\ X\cup Y\in{\sf FINITE}).$
Base case: $\emptyset\cup Y=Y$ is finite.
Induction step: Suppose $X=Z\cup\{b\}$ with $b\not\in Z$ and $Y\cap(Z\cup\{b\})=\emptyset$ and $Z$ finite. Then
@@
$X\cup Y$ $=$ $Z\cup(Y\cup\{b\})$ (6)
Since $Y\cap(Z\cup\{b\})=\emptyset$, $b\not\in Y$. Then by the definition of $\sf FINITE$}, $Y\cup\{b\}$ is finite. We have
$Z\cap(Y\cup\{b\})=Y\cap(Z\cup\{b\})=\emptyset.$
Then by the induction hypothesis, $Z\cup(Y\cup\{b\})$ is finite. Then by (6), $X\cup Y$ is finite. That completes the induction step. ∎
@@
Proof.
Suppose $x\sim y$. Then there exists $f:x\to y$ with $f$ one-to-one and onto. By Lemma 2.9, $f^{-1}:y\to x$ is a one-to-one function. Then we have for $u,v\in y$,
$u=v{\ \leftrightarrow\ }f^{1}(u)=f^{-1}(v).$ (7)
@@
Therefore $y$ has decidable equality. ∎
@@
Proof.
Let $q=f(c)$. Then $g:z\to y-\{q\}$. Suppose $g(u)=g(v)$. Then $f(u)=f(v)$. Since $f$ is one-to-one, $u=v$. Hence $g$ is one-to-one. Suppose $v\in y-\{q\}$. Since $f$ is onto, $v=f(u)$ for some $u\in z\cup\{c\}$; but $u\neq c$ since if $u=c$ then $v=f(u)=q$, but $v\neq q$ since $v\in y-\{q\}$. Then $u\in z$. Hence $g$ is onto. ∎
@@
Proof.
We prove by induction on finite sets $x$ that
@@
Base case: When $x=\emptyset$. Suppose $y\sim\emptyset$. Then $y=\emptyset$, so $y\in{\sf FINITE}$. That completes the base case.
@@
Therefore $y\in{\sf FINITE}$. That completes the induction step. ∎
@@
$\mathcal{P}(X)=\{Y:Y\subset X\}.$
We shall not make use of $\mathcal{P}(X)$, because there are “too many” subclasses of $X$. Consider, by contrast, the separable subclasses of $X$:
@@
$\mathcal{P}_{s}(X):=\{u:u\subseteq X\ \land\ X=u\cup(X-u)\}$
That is, $u$ is a separable subclass (or subset, which is synonymous) of $X$ if and only if $\forall y\in X\,(y\in u\ \lor\ y\not\in u)$. Classically, of course, every subset is separable, so we have $\mathcal{P}_{s}(X)=\mathcal{P}(X)$, but that is not something we can assert constructively. The formula in the definition is stratified, so the definition can be given in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. When working with finite sets, $\mathcal{P}_{s}(X)$ is a good constructive substitute for $SC(X)$. We illustrate this by proving some facts about $SC(X)$, before returning to the question of the proper constructive substitute for $SC(X)$ when $X$ is not necessarily finite.
@@
Let $x$ be a finite set. Then $\mathcal{P}_{s}(x)$ is also a finite set.
Remark. We cannot prove this with $SC(x)$ in place of $\mathcal{P}_{s}(x)$.
Proof.
The formula to be proved is
$x\in{\sf FINITE}\ {\rightarrow}\ \mathcal{P}_{s}(x)\in{\sf FINITE}.$
@@
Base case: $\mathcal{P}_{s}(\emptyset)=\{\emptyset\}$ is finite.
@@
By the induction hypothesis, $\mathcal{P}_{s}(x)\in{\sf FINITE}$. I say that the map $u\mapsto u\cup\{c\}$ is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$:
$f:=\{\langle u,y\rangle:u\in\mathcal{P}_{s}(x)\ \land\ y=u\cup\{c\}.\}$
The formula can be stratified by giving $c$ index 0, $u$ and $y$ index 1, $\mathcal{P}_{s}(x)$ index 2; then $\langle u,y\rangle$ has index 3 and we can give $f$ index 4. Hence $f$ is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$ as claimed. $f$ is a function since $y$ is uniquely determined as $u\cup\{c\}$ when $u$ is given. Also $f$ is one-to-one, since if $u\subseteq x$ and $v\subseteq x$ and $c\not\in x$, and $u\cup\{c\}=v\cup\{c\}$, then $u=v$. Define
@@
$A=\{u\cup\{c\}:u\in\mathcal{P}_{s}(x)\}.$ (8)
Then $\mathcal{P}_{s}(x)\sim A$, because $f:\mathcal{P}_{s}(x)\to A$ is one-to-one and onto. Since $\mathcal{P}_{s}(x)$ is finite (by the induction hypothesis), by Lemma 3.3, $\mathcal{P}_{s}(x)$ has decidable equality. Then $A$ has decidable equality, by Lemma 3.12. Since $A$ has decidable equality, and is similar to the finite set $\mathcal{P}_{s}(x)$, $A$ is finite, by Lemma 3.14.
@@
Left-to-right: Let $v\in\mathcal{P}_{s}(x\cup\{c\})$. Then $v$ is a separable subset of $x\cup\{c\}$. Then $c\in v\ \lor\ c\not\in v$. If $c\not\in v$ then $v\in\mathcal{P}_{s}(x)$. If $c\in v$
@@
We have $v-\{c\}\in\mathcal{P}_{s}(x)$, since $v\subset x\cup\{c\}$ and $v$ has decidable equality. Then $f(v-\{c\})\in Range(f)=A$. But $f(v-\{c\})=(v-\{c\})\cup\{c\}=v$. Therefore $v\in A$. Therefore $v\in A\cup\mathcal{P}_{s}(x)$, as desired. That completes the proof of the left-to-right direction of (9).
Right-to-left. Let $v\in A\cup\mathcal{P}_{s}(x)$. Then $v\in A\ \lor\ v\in\mathcal{P}_{s}(x)$.
Case 1, $v\in A$. Then by (8), $v$ has the form $v=u\cup\{c\}$ for some $u\in\mathcal{P}_{s}(x)$. Then $u\cup\{c\}\in\mathcal{P}_{s}(x\cup\{c\})$ as required.
Case 2, $v\in\mathcal{P}_{s}(x)$. First we note that if $c\not\in x$ then
$\mathcal{P}_{s}(x)\subseteq\mathcal{P}_{s}(x\cup\{c\})$
Therefore, since $v\in\mathcal{P}_{s}(x)$, we have $v\in\mathcal{P}_{s}(x\cup\{c\})$. That completes the proof of (9).
Note that $A$ and $\mathcal{P}_{s}(x)$ are disjoint, since every member of $A$ contains $c$, and no member of $\mathcal{P}_{s}(x)$ contains $c$, since $c\not\in x$. Then by Lemma 3.11 and (9), $\mathcal{P}_{s}(x\cup\{c\})\in{\sf FINITE}$, as desired. ∎
@@
Proof.
Let $a\in{\sf FINITE}$. By induction on finite sets $b$ we prove
$b\in{\sf FINITE}\ {\rightarrow}\ b\subseteq a\ {\rightarrow}\ a=(a-b)\cup b.$ (10)
@@
Base case: Suppose $b=\emptyset$. Then $b\subseteq a$, so we have to prove $a=(a-\emptyset)\cup\emptyset$, which is immediate. That completes the base case.
@@
We must show $x\in a$. If $x\in(a-(b\cup\{c\})$ then $x\in a$. If $x\in(b\cup\{c\})$ then $x\in a$, since by hypothesis $b\cup\{c\}\subseteq a$. That completes the right-to-left direction. That completes the induction step. ∎
@@
Proof.
By induction on finite sets $X$. When $X$ is the empty set, every subset of $X$ is the empty set, so every subset of $X$ is empty, and hence finite. Now let $X=Y\cup\{a\}$ with $a\not\in Y$ and $Y$ finite, and let $U$ be a separable subset of $X$; that is,
$\forall z\in X\,(z\in U\ \lor\ z\not\in U).$ (15)
@@
$\forall z\in Y\,(z\in V\ \lor\ z\not\in V)$ (16)
[…]
@@
Let $a$ and $b$ be finite sets with $b\subseteq a$. Then aba-b is also a finite set.
Proof.
We first prove the special case when $b$ is a singleton, $b=\{c\}$. That is,
$a\in{\sf FINITE}\ \land\ c\in a\ {\rightarrow}\ a-\{c\}\in{\sf FINITE}$ (17)
By Lemma 3.3, $a$ has decidable equality. Hence $a-\{c\}$ is a separable subset of $a$. Then by Lemma 3.19, it is finite. That completes the proof of (17).
@@
Base case: $\emptyset-b=\emptyset$ is finite.
@@
Since $b$ is finite, also $b-\{c\}$ is finite, by (17). Since $b-\{c\}\subseteq p$, by the induction hypothesis we have
@@
Case 2: $c\not\in b$. Then $b\subseteq p$, so by the induction hypothesis pbp-b is finite.
@@
Therefore aba-b is finite. That completes Case 2. That completes the induction step. ∎
@@
Proof.
The formula to be proved is stratified, with $\sf FINITE$} as a parameter, giving $u$ and $z$ index 0, $B$ index 1, and $R$ index 3. Therefore it is legal to prove it by induction on finite sets $B$.
Base case: $B=\emptyset$. Then $X\times Y=\emptyset$, so $R=\emptyset$. Then $\forall z\,\neg\,\exists u\in B\,\langle u,z\rangle\in R$, and therefore
@@
$\forall z\in X\,(\exists u\in B\,\langle u,z\rangle\in R$ ${\ \leftrightarrow\ }$ $(\exists u\in A\,\langle u,z\rangle\in R)\ \lor\ \langle c,z\rangle\in R)$ (18)
@@
$(\exists u\in B\,\langle u,z\rangle\in R)\ \lor\ \neg\,(\exists u\in B\,\langle u,z\rangle\in R)$ (19)
@@
${\ \leftrightarrow\ }$ $(\exists u\in A\,\langle u,z\rangle\in R\ \lor\ \langle c,z\rangle\in R)\ \lor\ (\neg\,(\exists u\in A\,\langle u,z\rangle\in R)\ \land\ \langle c,z\rangle\not\in R))$
${\ \leftrightarrow\ }$ $(\exists u\in A\,\langle u,z\rangle\in R\ \lor\neg\,\exists u\in A\,\langle u,z\rangle\in R\ \lor\langle c,z\rangle\in R)$
@@
But by the induction hypothesis, that holds. That completes the induction step. ∎
@@
Proof.
Since $b\in U$ and $c\not\in U$, we have $b\neq c$. Define $f:U\to Y$ by
@@
Since $X$ has decidable equality, $f$ is well-defined on $X$, and from the definitions of $f$ and $Y$ we see that $f:U\to Y$ and $f$ is onto. Ad one-to-one: suppose $f(u)=f(v)$. Since $X$ has decidable equality, $u$ and $v$ are either equal or not. If $u=v$, we are done. If $u\neq v$ then exactly one of $u,v$ is equal to $b$, say $u=b$ and $v\neq b$. Then xf(u)=cf(u)=c and $f(v)=v$. Since $f(u)=f(v)$ we have $v=c$. But $v\in U$ and $c\not\in U$, contradiction. ∎
@@
The class $X$ is infinite if $X\sim Y$ for some $Y\subseteq X$ with $Y\neq X$. [^10] Alternate definitions one might consider: $X$ is infinite if there is a similarity from $X$ to a subset of $X$ that omits a value; $X$ is infinite if XAX-A is inhabited, for every finite set $A$. Whether Dedekind infinite implies these properties is not known.
@@
Proof.
It suffices to show that every finite set is not infinite. The formula to be proved is